OpenStudy (anonymous):

solve the quadratic equation. x^2-3x=4

OpenStudy (shamil98):

x^2 - 3x - 4 = 0 factor it.

OpenStudy (swissgirl):

We first need to set it up in the form \(ax^2+bx+c=0\)

OpenStudy (swissgirl):

Well ya like @shamil98 has shown we subtract 4 from both sides to achieve this

OpenStudy (anonymous):

Ok I got that. I have x^2-3x-4=0 but I am not sure where to go from there.

OpenStudy (shamil98):

you can use the quadratic formula or just factor it to (x+b)(x-b) form. and then solve for (x+b)=0 and (x-b) = 0 to get your solutions

OpenStudy (swissgirl):

We need to find 2 numbers that when you add them they equal -3 and when you multiply them they equal -4

OpenStudy (anonymous):

Well that would be -4 and 1.

OpenStudy (shamil98):

correct.

OpenStudy (swissgirl):

Correct

OpenStudy (shamil98):

so, in (x+b)(x-b)=0 form it would be?

OpenStudy (swissgirl):

So basically then we replace -3x with -4x +1x

OpenStudy (swissgirl):

So our equation would look as follows \(x^2-4x+1x-4\)

OpenStudy (shamil98):

(x-4)(x+1) = 0 after you put it in that form x - 4 = 0 and x + 1 = 0 solve for x to get your solutions.

OpenStudy (swissgirl):

Then we group the first 2 terms and teh last 2 terms \( (x^2-4x)+(x-4)=0

OpenStudy (swissgirl):

\( (x^2-4x)+(x-4)=0\)

OpenStudy (swissgirl):

Then we factor out from each bracket x(x-4)+1(x-4)=0

OpenStudy (shamil98):

You made a mistake I believe.. @swissgirl

OpenStudy (swissgirl):

At which point?

OpenStudy (shamil98):

Well, she is asking how to find the solution, I have no idea what you are showing.. the original equation is x^2 - 3x - 4. factored out (x-4)(x+1)= 0 x = -1 x =4

OpenStudy (shamil98):

x^2 - 4x + x -4 can just be simplified to x^2 -3x - 4 .....

OpenStudy (swissgirl):

Yes but I am just demonstrating how you factored it out This is the typical way it is taught in school and when it becomes more challenging then it is imperative that you write it out in this fashion

OpenStudy (swissgirl):

@Katelynmorris which method are you familiar with?

OpenStudy (shamil98):

It looks more complicated with your method x^2 -3x -4 you know that the values are 1 and - 4. just put that like this (x-4)(x+1) = 0

OpenStudy (anonymous):

I am lost after the x^2-4x+1x-4=0..Solving by factoring

OpenStudy (swissgirl):

But @shamil98 it works this time but what happens if the equation was \(3x^2-3x-4=0\) ?

OpenStudy (anonymous):

Ok so after the (x-4)(x+1)=0 that gives me x how?

OpenStudy (anonymous):

listen to sham, he got da plan @_@

OpenStudy (shamil98):

(x-4)(x+1)=0 you put x-4=0 and x+1=0 to get your solutions .

OpenStudy (anonymous):

@Katelynmorris just look at daht smart score of his @_@ YO SWISS, DONT BE CHEESY <3 I loe u, but let L do his Justice <3

OpenStudy (shamil98):

3x2−3x−4=0 swiss, i would use the quadratic formula. instead of trying to factor it.

OpenStudy (swissgirl):

You can use whtvr it makes no diff as long as you get your answer

OpenStudy (shamil98):

Your method is much more complicated than it needs to be, just make it simple..

OpenStudy (swissgirl):

The only reason why I was showing him that method since that is the method taught in schools

OpenStudy (swissgirl):

Plus your method wont work for complicated equations trust me kid

OpenStudy (anonymous):

Ok so x= 4, -1?

OpenStudy (shamil98):

Yes, those are your solutions.

OpenStudy (anonymous):

Thanks.

OpenStudy (anonymous):

as a drunk person on peaches, let me tell u somethin swiss cheese, sham aint no kid, he be Justice @_@

OpenStudy (anonymous):

don't make me get my peaches loe <3 @Katelynmorris listen to sham, hes been doin dis for years @_@ just look at daht bronze 76 he got @_@

OpenStudy (anonymous):

every time you type @_@ you tag whoever that is lol

OpenStudy (anonymous):

LAWL XD

OpenStudy (anonymous):

OMG THEY ACTUALLY EXCIST o-o

OpenStudy (anonymous):

I have been using that face for so many times like this.........they will witness the spam of there lives when they come back XD

OpenStudy (anonymous):

stalking profile now

OpenStudy (anonymous):

XD